Please help me with this question. I'm getting really confused with the 3x. I' not sure where to start.
Find all of the exact solutions to the following equation. (Use the parameter k as necessary to represent any integer.)
cos(3x) =1/2

Answer by lwsshak3(11628)

Find all of the exact solutions to the following equation. (Use the parameter k as necessary to represent any integer.)
cos(3x) =1/2
3x=π/3+2πk, 5π/3+2πk, k=any integer
x=π/9+2πk, 5π/9+2πk, k=any integer